OpenStudy (anonymous):

Can someone help me with L'Hospital's Rule consisting with limits? what is the limit of [x/(x-1)]-[1/lnx] as x approaches 1? Thank you!

OpenStudy (kirbykirby):

Ok I found a solution to this. It's a bit long to type so bear with me

OpenStudy (anonymous):

okay! :) thank you!

OpenStudy (kirbykirby):

HR in the following is for L'Hôpital's Rule \[\large \lim_{x\rightarrow 1 }\left( \frac{x}{x-1}-\frac{1}{\ln x}\right)\\ =\large \lim_{x\rightarrow 1}\left( \frac{x(\ln x)}{(x-1)(\ln x)}-\frac{(x-1)}{\ln x(x-1)}\right), \text{put same denominator}\\ =\large \lim_{x\rightarrow 1}\left( \frac{x\ln x -x +1}{(x-1)\ln x}\right),\text{form }\frac{0}{0}\implies \text{use l'HR}\\ =\large \lim_{x\rightarrow 1}\left( \frac{\ln x+x\frac{1}{x}-1}{\ln x+(x-1)\frac{1}{x}}\right)\\ =\large \lim_{x\rightarrow 1}\left( \frac{\ln x}{\ln x+1-\frac{1}{x}}\right)\text{form }\frac{0}{0}\implies \text{use l'HR}\\ =\large \lim_{x\rightarrow 1}\left( \frac{\frac{1}{x}}{\frac{1}{x}+\frac{1}{x^2}}\right)\\ =\frac{\frac{1}{1}}{\frac{1}{1}+\frac{1}{1^2}}\\ =\frac{1}{2}\]

OpenStudy (anonymous):

Okay, I see where I made my mistake now. I didn't take the derivative properly. Thank you!!!! You're awesome.

OpenStudy (kirbykirby):

:)